An initial capital of $20,000 was put uo for a new project that will produce an annual income of $7500 for 7 years and then will have a salvage value of $4000 at that time. Annual expenses for its operation and maintenance amounts to $3000. If money is worth 10% compounded annually, determine if this project is profitable using External rate of return method.

PLEASE DO NOT USE COMPUTER PROGRAMS (Excel, Matlab)
Answer #1

ANSWER:

I = 10%

Fv of positive cash flows = annual income(f/a,i,n) + salvage value = 7,500(f/a,10%,7) + 4,000 = 7,500 * 9.487 + 4,000 = 71,152.5 + 4,000 = 75,152.5

pv of negative cash flows = initial capital + o and m amount(p/a,i,n) = 20,000 + 3,000(p/a,10%,7) = 20,000 + 3,000 * 4.868 = 20,000 + 14,604 = 34,604

n = 7

err = (fv of positive cash flows / pv of negative cash flows) ^ 1 / n - 1

err = (75,152.5 / 34,604) ^ 1 / 7 - 1

err = (2.1717) ^ 1 / 7 - 1

err = 1.1171 - 1

err = 0.1171 or 11.71%

so the external rate of return is 11.71%

since the err > marr that is 11.71% > 10%

therefore we will accept the project as it will be profitable.